please help solve this problem 18xSQUARE+50=60X
Answer by checkley71(8403)

18X^2+50=-60X
16X^2-60X+50=0
(8X-10)(2X-5)=0
8X-10=0
8X=10
X=10/8
X=5/4 ANSWER.
2X=5
2X=5
X=5/2 ANSWER.